# AfriDoc-MT: Document-level MT Corpus for African Languages

### v1

@@ Abstract
This paper introduces AfriDoc-MT, a document-level multi-parallel translation dataset covering English and five African languages: Amharic, Hausa, Swahili, Yorùbá, and Zulu. The dataset comprises 334 health and 271 information technology news documents, all human-translated from English to these languages. We conduct document-level translation benchmark experiments by evaluating neural machine translation (NMT) models and large language models (LLMs) for translations between English and these languages, at both the sentence and pseudo-document levels. These outputs are realigned to form complete documents for evaluation. Our results indicate that NLLB-200 achieved the best average performance among the standard NMT models, while GPT-4o outperformed general-purpose LLMs. Fine-tuning selected models led to substantial performance gains, but models trained on sentences struggled to generalize effectively to longer documents. Furthermore, our analysis reveals that some LLMs exhibit issues such as under-generation, repetition of words or phrases, and off-target translations, especially for African languages.
@@ 1 Introduction
The field of machine translation (MT) has seen notable progress in the past years, particularly with neural machine translation (NMT) models achieving close to human performance in many high-resource languages Vaswani et al. (2017); Akhbardeh et al. (2021); Mohammadshahi et al. (2022); Team et al. (2024); Yuan et al. (2023); Kocmi et al. (2023). However, efforts have primarily been concentrated on sentence-level translation, without the use of inter-sentential context.
@@
In recent years, there has been interest in document-level translation (i.e. the holistic translation of multiple sentences), where sentences are translated with their context rather than in isolation. Document-level translation is important in order to capture discourse relations Bawden et al. (2018); Voita et al. (2018); Maruf et al. (2021), maintain consistency and coherence across sentences Herold and Ney (2023), particularly for technical domains, but poses unique challenges, such as how to handle longer documents Wang et al. (2024b) given the limited context size of translation models. Current efforts have primarily focused on high-resource language directions, where document-level datasets are readily available Lopes et al. (2020); Feng et al. (2022); Wu et al. (2023); Wang et al. (2023); Wu et al. (2024), and so far there has been no work on low-resource African languages. Developing and evaluating document-level MT systems for low-resource languages is a useful and under-studied direction, which requires the creation of datasets.
| Dataset | #Langs. | Multiway | Domain | Type | #Sents. |
| TICO-19 (Anastasopoulos et al., 2020) | ✓ | health | document-level | k | |
| MAFAND-MT (Adelani et al., 2022) | ✗ | news | sentence-level | k-k | |
| FLORES-200 (NLLB Team et al., 2022) | ✓ | general | sentence-level | k | |
| NTREX-128 Federmann et al. (2022) | ✓ | news | sentence-level | k | |
| AfriDoc-MT (Ours) | ✓ | tech, health | document-level | k |
@@
To fill this gap, we present AfriDoc-MT, a document-level translation dataset for English from and into five African languages: Amharic, Hausa, Swahili, Yorùbá, and Zulu, created through the manual translation of English documents. The dataset, drawn from 334 health documents and 271 tech documents, contains sentences per domain for each language pair. In addition, AfriDoc-MT supports multi-way translation, allowing translations not only between English and the African languages but also between any two of the languages covered.
@@
We conduct a comprehensive set of document translation benchmark experiments on AfriDoc-MT, using sentence-level and pseudo-document translation due to most models’ limited context length, and then realigning them to form complete documents. We evaluate performance using automatic metrics and compare the results of encoder-decoder models with decoder-only LLMs across both domains. Our results demonstrate that NLLB-200, both before and after fine-tuning on AfriDoc-MT, excels in sentence translation, surpassing all other models. GPT-4o performs equally well for sentences and pseudo-documents, while other decoder-only models lag behind. In addition, we use GPT-4o as a proxy for human evaluation to compare documents translated sentence by sentence with those translated as pseudo-documents. The evaluation shows that, on average, pseudo-document translations are more fluent and have fewer errors than sentence-level translations. We conducted additional analyses on the models outputs to better understand their behavior. Our analyses show that LLMs often undergenerate, generate repetitions, and produce off-target translations, especially when translating into African languages.
@@ 2 Related Work
Several MT datasets exist for African languages, including web-mined datasets such as WikiMatrix Schwenk et al. (2021a) and CCMatrix Schwenk et al. (2021b). However, they have been adjudged to be of poor quality for certain low-resource subsets, including African languages Kreutzer et al. (2022). There are also well curated datasets for African languages including the Bible McCarthy et al. (2020), JW300 Agić and Vulić (2019)111The dataset is no longer available for use. and MAFAND-MT Adelani et al. (2022), which are from religious and news domains.
There exist several MT evaluation benchmark datasets for African languages. They can be categorized into two kinds. First, evaluation datasets specifically designed for translating into or from African languages (Ezeani et al., 2020; Azunre et al., 2021; Adelani et al., 2021, 2022, inter alia). Second, benchmark datasets covering many languages, including African languages. For example, TICO-19 (Anastasopoulos et al., 2020), NTREX-128 Federmann et al. (2022), FLORES-101 Goyal et al. (2022) and FLORES-200 Team et al. (2024) are a few such datasets. However, most of these datasets are designed for sentence-level MT, primarily drawn from religious or news domains, although some consist of translated sentences originating from the same document. To the best of our knowledge, only TICO-19, a health domain translation benchmark, has the potential to be used for document-level MT, while it is restricted to topics related to COVID-19. Table 1 gives a comparison of the most related existing benchmarks.
@@
Document-level NMT aims to overcome the limitations of sentence-level systems by translating an entire document as a whole. Similar to context-aware NMT, which involves translating segments with additional, localized context, it differs in that it involves in principle translating an entire document holistically. Both document-level and context-aware MT allow for the possibility of improving translation quality for context-dependent phenomena such as coreference resolution Müller et al. (2018); Bawden et al. (2018); Voita et al. (2018); Herold and Ney (2023), lexical disambiguation Rios Gonzales et al. (2017); Martínez Garcia et al. (2019), and lexical cohesion Wong and Kit (2012); Garcia et al. (2014, 2017); Bawden et al. (2018); Voita et al. (2019). Various methods have been proposed to extend sentence-level models to capture document-level context Tiedemann and Scherrer (2017); Libovický and Helcl (2017); Bawden et al. (2018); Miculicich et al. (2018); Sun et al. (2022). The emergence of LLMs, such as GPT-3 Brown et al. (2020), Llama Dubey et al. (2024) and Gemma Gemma Team et al. (2024), has transformed NLP, including for MT Zhu et al. (2024a, c); Lu et al. (2024). Pre-trained on vast amounts of text, LLMs can effectively manage long-range dependencies, making them in principle well-suited for document-level translation. While these models have shown promising results for high-resource languages Wu et al. (2023); Wang et al. (2023); Wu et al. (2024), research remains limited for low-resource languages Ul Haq et al. (2020).
@@ 3 AfriDoc-MT Corpus
We cover five languages from the two most common African language families: Afro-Asiatic and Niger-Congo. Three languages belong to the Niger-Congo family: Swahili (North-East Bantu), Yorùbá (Volta-Niger) and isiZulu (Southern Bantu). The other two languages belong to the Afro-Asiatic family: Amharic (Semitic) and Hausa (Chadic). Some details of these languages are given in Table 2. Each of these languages has over million speakers and is spoken in different regions of Sub-Saharan Africa: East (Amharic, Swahili), West (Hausa, Yorùbá), and South (isiZulu). All of them use the Latin script except for Amharic, which uses the Ge’ez script. The Latin-script languages use the Latin alphabet with the omission of some letters and the addition of new ones, and the use of diacritics (e.g., Yorùbá). The languages are tonal, except for Amharic and Swahili. Just like English, all languages follow the subject-verb-object word order. Refer to Adelani (2022) for a comprehensive overview of the characteristics of these languages.
@@
We scraped English articles from the websites of Techpoint Africa222https://techpoint.africa/ and the World Health Organization (WHO)333https://www.who.int/health-topics,444https://www.who.int/news-room/. The articles cover different topics of different lengths with an average length of 30 and 37 sentences for health and tech respectively. While our corpus is initially structured at the article level, we aim to make it suitable for sentence-level translation tasks as well. To achieve this, we segmented the raw articles into sentences using NLTK Bird et al. (2009). To ensure high segmentation quality, we recruited a linguist and a professional translator to verify the correctness of the segmentation and made corrections as needed. Finally, we selected 10k sentences from the collected texts which resulted in 334 and 271 English articles from the health and tech domains respectively.
@@
We translated the extracted 10k English sentences to the 5 African languages through 4 expert translators per language.555Each translator was paid for sentences. The translators were recruited through a language coordinator who is also a native speaker of the language. The 10k sentences were distributed equally among the translators and the translations were done in-context (i.e. the translators translated on the sentence level but had access to the the whole document). Due to the domain-specific nature of the task, before starting the translation process, we conducted a translation workshop, during which three translation experts shared their experiences in creating terminologies and they also shared existing resources with the translators including a short translation guideline (Section A.1).
@@
Quality control was conducted using automated quality estimation, followed by manual inspections by our language coordinators. Due to the volume of translations, we also used automated methods to verify the quality of translations, including language identification to confirm translations were in the target language using Google’s DETECTLANG function. We used Quality Estimation (QE), specifically AfriCOMET (Wang et al., 2024a), to assess translation quality. Translations scoring below 0.65 were jointly reviewed by translators and language coordinators (see Section A.2).
@@
We created train, development (dev), and test splits for each domain. To prevent data leakage, we first identified documents that shared sentences with the same English translation and assigned these documents to the training set. The dev and test sets are derived from the remaining documents. The dev set comprises documents containing 800 to 1000 sentences, whereas the test set includes documents with 1800 to 2000 sentences. This approach ensures that our evaluation dataset sizes are comparable to other popular benchmark datasets such as FLORES and NTREX. Table 3 shows some data statistics, and we provide more data statistics in Appendix A666We will release AfriDoc-MT on https://github.com/masakhane-io/afridoc-mt..
@@ 4 Benchmark Experiments
Given the AfriDoc-MT data, we conducted both sentence- and document-level translation, evaluating two types of models: encoder-decoder and decoder-only models. While the majority of these models are open-source, we also evaluated two proprietary models of the same type. Our evaluation primarily focuses on document-level translation, reflecting the availability of our document-level translation corpus. For completeness, we also conduct a series of sentence-level experiments, with the results presented in Appendix C.
@@ 4.1 Models
We evaluate five kinds of open encoder-decoder model including Toucan Elmadany et al. (2024); Adebara et al. (2024), M2M-100 Fan et al. (2020), NLLB-200 Team et al. (2024), MADLAD-400 Kudugunta et al. (2023), and Aya-101 Üstün et al. (2024). Toucan is an Africa-centric multilingual MT model that supports 150 African language pairs. In comparison, M2M-100, NLLB-200, and MADLAD-400 are multilingual MT models that cover 100, 200, and 450 language pairs, respectively. Aya-101 is an instruction-tuned mT5 model Xue et al. (2021), covering languages that is capable of translating between different languages including the African languages considered in AfriDoc-MT.
@@
We also evaluate open and closed decoder-only models. The open models include LLama3.1 Dubey et al. (2024), Gemma2 Gemma Team et al. (2024), instruction-tuned versions of LLama3.1 and Gemma2, and LLaMAX3 Lu et al. (2024), which is a Llama3-based model with continued pre-training on over 100 languages including several African languages, whereas the other models are English-centric. The closed models we test are OpenAI GPT models (GPT-3.5 Turbo and GPT-4o) OpenAI (2024), which have been shown to have document-level translation ability Wang et al. (2023). Although the language coverage of most of these models is not well documented, they demonstrate some understanding of African languages Adelani et al. (2024b); Bayes et al. (2024), though not at the same level as English, which is the primary language in their training data.
We present the result of models in total, including the 1.2B version of Toucan, 1.3B and 3.3B versions of NLLB-200, 3B and 13B versions of MADLAD-400 and Aya-101 respectively. We also have the 8B instruction tuned version of LLama3.1 (LLama3.1-IT), 9B version of Gemma-2 (Gemma2-IT), and LLaMAX3-Alpaca777we refer to it as LLaMAX3-Alp in the results tables.. We provide more description of the models in Section B.1.
Supervised finetuning of the models
For sentence-level evaluation, we jointly fine-tune NLLB-200 with 1.3B parameters on the 30 language directions and on the two domains to make the models more specialized. Similarly, we did supervised fine-tuning on LLaMAX3 and LLama3.1 using the prompt augmentation approach from Zhu et al. (2024b), and shown in Section B.4. We chose these two models because LLaMAX3 is already adapted to several languages including our languages of interest, and LLama3.1 because of its long context window. We perform SFT on LLaMAX3 and LLama3.1 for document-level translation, using pseudo-documents with =10. We refer to each system as {model_name}-SFTk888we denote models finetuned on sentences as {model_name}-SFT or {model_name}-SFT1 .

### v2

@@ Abstract
This paper introduces AfriDoc-MT, a document-level multi-parallel translation dataset covering English and five African languages: Amharic, Hausa, Swahili, Yorùbá, and Zulu. The dataset comprises 334 health and 271 information technology news documents, all human-translated from English to these languages. We conduct document-level translation benchmark experiments by evaluating the ability of neural machine translation (NMT) models and large language models (LLMs) to translate between English and these languages, at both the sentence and pseudo-document levels, the outputs being realigned to form complete documents for evaluation. Our results indicate that NLLB-200 achieves the best average performance among the standard NMT models, while GPT-4o outperforms general-purpose LLMs. Fine-tuning selected models leads to substantial performance gains, but models trained on sentences struggle to generalize effectively to longer documents. Furthermore, our analysis reveals that some LLMs exhibit issues such as under-generation, over-generation, repetition of words and phrases, and off-target translations, specifically for translation into African languages.111The data can be accessed on HuggingFace and GitHub.
@@ 1 Introduction
The field of machine translation (MT) has seen notable progress in the past years, with neural machine translation (NMT) models achieving close to human performance in many high-resource language directions (Vaswani et al., 2017; Akhbardeh et al., 2021; Mohammadshahi et al., 2022; Yuan et al., 2023; Kocmi et al., 2023; NLLB Team et al., 2024). However, efforts have primarily been concentrated on sentence-level translation, without the use of inter-sentential context.
@@
In recent years, there has been interest in document-level translation (i.e. the holistic translation of multiple sentences), where sentences are translated with their context rather than in isolation. Document-level translation is important in order to capture discourse relations (Bawden et al., 2018; Voita et al., 2018; Maruf et al., 2021), maintain consistency and coherence across sentences (Herold and Ney, 2023), particularly for technical domains, but poses unique challenges, such as how to handle longer documents (Wang et al., 2024b) given the limited context size of translation models. Current efforts have primarily focused on high-resource language directions, where document-level datasets are readily available (Lopes et al., 2020; Feng et al., 2022; Wu et al., 2023; Wang et al., 2023; Wu et al., 2024), and so far there has been no work on low-resource African languages. Developing and evaluating document-level MT systems for low-resource languages is a useful and under-studied direction, which requires the creation of datasets.
| Dataset | #Langs. | Multiway | Domain | Type | #Sents. | #Docs. |
| TICO-19 (Anastasopoulos et al., 2020) | 12 | ✓ | health | document-level | 4k | 30 |
| MAFAND-MT (Adelani et al., 2022) | 16 | ✗ | news | sentence-level | 4k-35k | - |
| FLORES-200 (NLLB Team et al., 2022) | ✓ | general | sentence-level | 3k | - | |
| NTREX-128 Federmann et al. (2022) | 24 | ✓ | news | sentence-level | 1.9k | - |
| AfriDoc-MT (Ours) | 5 | ✓ | tech, health | document-level | 10k | 271-334 |
@@
To fill this gap, we present AfriDoc-MT, a document-level translation dataset for English from and into five African languages: Amharic, Hausa, Swahili, Yorùbá, and Zulu, created through the manual translation of English documents. It consists of 334 health documents and 271 tech documents. In addition, AfriDoc-MT supports multi-way translation, allowing translations not only between English and the African languages but also between any two of the languages covered.
@@
We conduct a comprehensive set of document translation benchmark experiments on AfriDoc-MT, using sentence-level and pseudo-document translation due to most models’ limited context length, and then realigning them to form complete documents. We evaluate performance using automatic metrics and compare the results of encoder-decoder models with decoder-only LLMs across both domains. Our results demonstrate that NLLB-200, both before and after fine-tuning on AfriDoc-MT, excels in sentence translation, surpassing all other models. GPT-4o performs equally well for sentences and pseudo-documents, while other decoder-only models lag behind. In addition to automatic metrics, we use GPT-4o as a judge, human evaluation, and qualitative assessment to compare documents translation carried out sentence-by-sentence and as pseudo-documents for selected models. The evaluation shows that GPT-4o is generally unreliable for assessing document translations into African languages. However, we observe agreement between other evaluation methods, all indicating that sentence-by-sentence translation results in better document-level translation into African languages. We conduct additional analyses of the models’ outputs to better understand their behavior and why they under-perform when translating pseudo-documents. They show that LLMs often under-generate, contain repetitions, and produce off-target translations, especially when translating into African languages.
@@ 2 Related Work
Several MT datasets exist for African languages, including web-mined datasets such as WikiMatrix (Schwenk et al., 2021a) and CCMatrix (Schwenk et al., 2021b). However, they have been adjudged to be of poor quality for certain low-resource subsets, including African languages (Kreutzer et al., 2022). There are also well curated datasets for African languages including the Bible (McCarthy et al., 2020), JW300 (Agić and Vulić, 2019)222The dataset is no longer available for use. and MAFAND-MT (Adelani et al., 2022), which are from religious and news domains.
There exist several MT evaluation benchmark datasets for African languages. They can be categorized into two kinds. First, evaluation datasets specifically designed for translating into or from African languages (Ezeani et al., 2020; Azunre et al., 2021; Adelani et al., 2021, 2022, inter alia). Second, benchmark datasets covering many languages, including African languages. For example, TICO-19 (Anastasopoulos et al., 2020), NTREX-128 (Federmann et al., 2022), FLORES-101 (Goyal et al., 2022) and FLORES-200 (NLLB Team et al., 2024) are a few such datasets. However, most of these datasets are designed for sentence-level MT, primarily drawn from religious or news domains, although some consist of translated sentences originating from the same document. To the best of our knowledge, only TICO-19, a health domain translation benchmark, has the potential to be used for document-level MT, while it is restricted to topics related to COVID-19. Table˜1 gives a comparison of the most relevant existing benchmarks.
@@
Document-level NMT aims to overcome the limitations of sentence-level systems by translating an entire document as a whole. Similar to context-aware NMT, which involves translating segments with additional, localized context, it differs in that it involves in principle translating an entire document holistically. Both document-level and context-aware MT allow for the possibility of improving translation quality for context-dependent phenomena such as coreference resolution (Müller et al., 2018; Bawden et al., 2018; Voita et al., 2018; Herold and Ney, 2023), lexical disambiguation (Rios Gonzales et al., 2017; Martínez Garcia et al., 2019), and lexical cohesion (Wong and Kit, 2012; Garcia et al., 2014, 2017; Bawden et al., 2018; Voita et al., 2019). Various methods have been proposed to extend sentence-level models to capture document-level context (Tiedemann and Scherrer, 2017; Libovický and Helcl, 2017; Bawden et al., 2018; Miculicich et al., 2018; Sun et al., 2022). The emergence of LLMs, such as GPT-3 (Brown et al., 2020), Llama (Dubey et al., 2024) and Gemma (Gemma Team et al., 2024), has transformed NLP, including for MT (Zhu et al., 2024a, c; Lu et al., 2024). Pre-trained on vast amounts of text, LLMs can effectively manage long-range dependencies, making them in principle well-suited for document-level translation. While these models have shown promising results for high-resource languages (Wu et al., 2023; Wang et al., 2023; Wu et al., 2024), research remains limited for low-resource languages (Ul Haq et al., 2020).
@@ 3 AfriDoc-MT Corpus
We cover five languages from the two most common African language families: Afro-Asiatic and Niger-Congo. Three languages belong to the Niger-Congo family: Swahili (North-East Bantu), Yorùbá (Volta-Niger) and isiZulu (Southern Bantu). The other two languages belong to the Afro-Asiatic family: Amharic (Semitic) and Hausa (Chadic). The choice of languages was based on geographical representation, speaking population, and web coverage (which we consider as a proxy for the potential performance of existing models on these languages). Furthermore, each of these languages has over million speakers. All of them use the Latin script except for Amharic, which uses the Ge’ez script. The Latin-script languages use the Latin alphabet with the omission of some letters and the addition of new ones, and the use of diacritics (e.g., Yorùbá). The languages are tonal, except for Amharic and Swahili. Just like English, all languages follow the subject-verb-object word order. Refer to Adelani (2022) for a comprehensive overview of the characteristics of these languages. Table˜2 shows summarized details.
@@
We scraped English articles from the websites of Techpoint Africa333https://techpoint.africa/ and the World Health Organization (WHO).444https://www.who.int/health-topics,555https://www.who.int/news-room/ The articles cover different topics of different lengths with an average length of 30 and 37 sentences for health and tech respectively. While our corpus is initially structured at the article level, we aim to make it suitable for sentence-level translation tasks as well. To achieve this, we segmented the raw articles into sentences using NLTK (Bird et al., 2009). To ensure high segmentation quality, we recruited a linguist and a professional translator to verify the correctness of the segmentation and make corrections as needed. Finally, we selected 334 and 271 English articles/documents from the health and tech domains respectively, which represents 10k sentences each per domain.
@@
We translated the extracted 10k English sentences to the 5 African languages through 4 expert translators per language.666Each translator was paid for sentences. The translators were recruited through a language coordinator who is also a native speaker of the language. The 10k sentences were distributed equally among the translators and the translations were done in-context (i.e. the translators translated on the sentence level but had access to the whole document). Due to the domain-specific nature of the task, before starting the translation process, we conducted a translation workshop, during which three translation experts shared their experiences in creating terminologies and they also shared existing resources with the translators including short translation guidelines (Section˜A.1).
@@
Quality control was conducted using automated quality estimation, followed by manual inspections by our language coordinators. We also used Quality Estimation (QE), specifically AfriCOMET (Wang et al., 2024a),777https://huggingface.co/masakhane/africomet-qe-stl Given a translated sentence in any African language and its corresponding source English sentence, AfriCOMET generates a score between 0 and 1, where 0 indicates poor quality and higher values signify better quality. The translators, in collaboration with the language coordinators, were tasked with reviewing instances that had quality estimation scores below 0.65. This step was essential to identify and correct low-quality translations.
Figure˜3 shows the distribution of final quality scores for five languages across both domains. Manual inspection indicates that QE scores below 0.65 do not necessarily reflect poor translations, consistent with Adelani et al. (2024b), likely due to domain shift, translation length, and other factors, which warrant further investigation.
@@
We created train, development (dev), and test splits for each domain. To prevent data leakage, documents sharing sentences with the same English translation were assigned to the training set. The dev set contains 25–33 documents, and the test set 59–61 documents, drawn from the remaining data. Table˜4 shows the average number of whitespace-separated tokens per sentence across domains and languages, including English. The health domain has more tokens than tech. Hausa and Yorùbá are longer than English, likely due to their descriptive nature, while Swahili is similar in length. Amharic and Zulu are relatively shorter, reflecting interesting linguistic properties. Table˜3 provides additional data statistics. Table˜3 shows some data statistics.
The health data is licensed under CC BY-NC-SA 3.0, while the tech data is licensed under CC BY-NC-SA 4.0.
@@ 4 Benchmark Experiments
Given the AfriDoc-MT data, we conducted both sentence- and document-level translation, evaluating two types of models: encoder-decoder and decoder-only models. While the majority of these models are open-source, we also evaluated two proprietary models of the same type. Our evaluation primarily focuses on document-level translation, reflecting the availability of our document-level translation corpus. For completeness, we also conduct a series of sentence-level experiments, with the results presented in Appendix˜C.
| Domain | eng | amh | hau | swa | yor | zul |
| Sentence | ||||||
| health | ||||||
| tech | ||||||
| Document | ||||||
| health | ||||||
| tech | ||||||
@@ 4.1 Models
We evaluate five kinds of open encoder-decoder model including Toucan (Elmadany et al., 2024; Adebara et al., 2024), M2M-100 (Fan et al., 2020), NLLB-200 (NLLB Team et al., 2024), MADLAD-400 (Kudugunta et al., 2023), and Aya-101 (Üstün et al., 2024). Toucan is an Afro-centric multilingual MT model supporting 150 African language pairs. In comparison, M2M-100, NLLB-200, and MADLAD-400 cover between 100 and 450 language pairs. Aya-101, an instruction-tuned mT5 model (Xue et al., 2021), supports 100 languages and can translate between various languages, including those considered in AfriDoc-MT.
@@
We also evaluate open and closed decoder-only models. Open models include LLama3.1 (Dubey et al., 2024), Gemma2 (Gemma Team et al., 2024), their instruction-tuned variants, and LLaMAX3 (Lu et al., 2024)—a LLama3-based model further pre-trained on 100+ languages, including several African ones. The closed models include OpenAI GPT models (GPT-3.5 Turbo and GPT-4o) (OpenAI, 2024), which have been shown to have document-level translation ability (Wang et al., 2023). While their language coverage is not well documented, they show some ability to handle African languages (Adelani et al., 2024b; Bayes et al., 2024), though far below their performance in English, their primary training language.
We present the result of models in total, including the 1.2B version of Toucan, 1.3B and 3.3B versions of NLLB-200, 3B and 13B versions of MADLAD-400 and Aya-101 respectively. We also have the 8B instruction tuned version of LLama3.1 (LLama3.1-IT), 9B version of Gemma-2 (Gemma2-IT), and LLaMAX3-Alpaca.888We refer to it as LLaMAX3-Alp in the results tables. We provide more description of the models in Section˜B.1.
Supervised fine-tuning of the models
For sentence-level evaluation, we jointly fine-tune NLLB-200 with 1.3B parameters on the 30 language directions and on the two domains to make the models more specialized. Similarly, we did supervised fine-tuning (SFT) on LLaMAX3 and LLama3.1 using the prompt augmentation approach from (Zhu et al., 2024b), as shown in Section˜B.4. We chose these two models because LLaMAX3 is already adapted to several languages including our languages of interest, and LLama3.1 because of its long context window. We perform SFT on LLaMAX3 and LLama3.1 for document-level translation, using pseudo-documents with =10. We refer to each system as {model_name}-SFTk.999We denote models finetuned on sentences as {model_name}-SFT or {model_name}-SFT1